OpenStudy (anonymous):

label all intercepts f(x)= -√ x-3 how do i get (7,-2) as an answer

OpenStudy (mertsj):

Replace x with 7

OpenStudy (mertsj):

\[f(x)=-\sqrt{7-3}=-\sqrt{4}=-2\]

OpenStudy (mertsj):

So if x is 7, f(x) is -2 resulting in the ordered pair (7,-2)

OpenStudy (anonymous):

how do i know thats an intercept

OpenStudy (mertsj):

It is NOT an intercept. An x intercept has a y value of 0 and a y intercept has an x value of 0

OpenStudy (mertsj):

You asked how to get (7,-2)

OpenStudy (mertsj):

It does not cross the y axis and the x intercept is (3,0)